# PhreshPhish: A Real-World, High-Quality, Large-Scale Phishing Website Dataset and Benchmark

### v1

@@ 1 Introduction
Phishing attacks are a pervasive and evolving threat that uses social engineering techniques to deceive millions of Internet users annually. According to the latest Phishing Activity Trends Report apwg-report-2024q4 released by the Anti-Phishing Working Group (APWG) apwg , almost a million phishing attacks were observed just in the fourth quarter of 2024. These attacks are a central component of modern cyber kill chains cyberkillchain , often involving campaigns that lure users to spoofed websites designed to trick them into revealing sensitive information, such as login credentials and financial data. The data thus obtained often serves as an initial entry point to fraud, identity theft and large-scale data breaches, resulting in substantial financial and reputational damage. In 2023 alone, phishing contributed significantly to the estimated \$12.5 Billion in cyberfraud losses reported to the FBI fbi-ic3report-2023 .
@@
Phishing remains a popular attack vector due to its effectiveness, low cost, and ease of deployment. A major contributing factor is the widespread availability of phishing kits cova2008there ; bijmans2021catching . These kits contain complete phishing websites in a ready-to-deploy package, and can often be downloaded for free or even purchased on the dark web markets dark-web-2018 , making such attacks accessible even to unskilled adversaries.
@@
Despite considerable research efforts, detecting phishing sites remains a challenging task das2019sok . Various approaches have been proposed to detect phishing, including use of blacklists, whitelists, heuristic-based techniques and machine learning methods. However, its adversarial nature, with attackers constantly adapting to avoid detection, and lack of good quality data undermines these defense approaches. Phishing websites are typically short-lived, highly dynamic, and often employ techniques such as obfuscation, cloaking 9519414 , and fast-flux hosting nagunwa2022machine .
[…]
@@
Figure 1: Our end-to-end pipeline consists of four distinct stages and we show the benign-to-phishing count at each step: (1) benign and phishing HTML is collected from the web using a real browser to ensure high fidelity; (2) the retrieved HTML is cleaned and assessed for quality using a combination of automated heuristics and human annotation; (3) train and test splits are created temporally and pruned to minimize leakage; (4) benchmark datasets are created by applying a set of diversity, difficulty and base rate filters.
@@
State-of-the-art machine learning-based phishing detection systems can be categorized into two broad approaches: (1) Direct phishing detection, where models are trained to classify webpages directly as phishing or benign whittaker2010large ; mohammad2015phishing ; shahrivari2020phishing ; aljofey2022effective . These models rely on feature extraction or representation learning from a (partially) labeled dataset of webpages. (2) Target brand-based phishing detection, where models are trained to identify the brand a webpage attempts to mimic, and phishing is inferred from inconsistencies between the webpage’s URL and the legitimate URLs associated with the identified brand afroz2011phishzoo ; lin2021phishpedia ; li2024phishintel . While the latter approach does not require a dataset of phishing webpages, it relies on a curated set of target brands, and typically incurs much higher detection latency.
@@
While the collection of large-scale web datasets is itself nontrivial, successfully scraping phishing pages–which behave adversarially to avoid detection–presents its own set of unique challenges beyond those typically encountered in general web scraping. In addition to the fact that the modern web is increasingly dynamic, issues such as ephemerality, cloaking, and staleness significantly complicate the collection of high-quality datasets which may be used to train and evaluate phishing detection models. These challenges necessitate the development of robust methodologies for collecting, cleaning, and validating phishing datasets to ensure their quality and utility for research purposes. We believe that these obstacles, discussed in greater detail in Section 3.1, contribute to the overall lack of high-quality datasets in this area. Although a search for publicly available phishing datasets would appear to yield promising results, we observed substantial quality issues with many datasets (see Appendix B). We also could not find any suitable benchmark datasets that facilitated a robust comparative evaluation of detection techniques despite the significant role that phishing plays in cybercrime.
@@
In this work, we present PhreshPhish, a large-scale, high-quality phishing website dataset collected over several months. To the best of our knowledge, this is the largest publicly available phishing website data and offers significantly higher quality compared to other such public datasets. Further, we create realistic benchmark datasets for comparison between detection methods. Figure 1 provides a schematic of the entire data pipeline from obtaining the phishing and benign URLs to finally producing the training, test and benchmark datasets.
To summarize, we make the following key contributions:
@@
  • We also release a suite of benchmark datasets to standardize the evaluation of detection methods. These benchmarks span 5 different base rates, ranging from 0.05% to 5%, making them more representative of real-world settings; further, we employ filters to increase dataset difficulty and diversity, and minimize data leakage.
@@
  • We evaluate and compare a number of baseline methods on our benchmark dataset to demonstrate the utility of our dataset and benchmarks. We also provide a detailed analysis of the performance of these methods, including their strengths and weaknesses, and discuss the implications for future research in this area.
2 Related Work
2.1 Phishing datasets
While an internet search for phishing website datasets returns a large number of hits, the actual number of datasets of a substantial size containing both URL and raw HTML data are very few. Most such datasets only contain website URLs, but no HTML (see Appendix B); others contain phishing emails or texts. A few datasets on Kaggle kaggle-huntingdata11 ; kaggle-haozhang1579 ; kaggle-jackcavar ; kaggle-aljofey and Hugging Face huggingface-ealvaradob ; huggingface-huynq3Cyradar contain raw HTML, however these are either smaller datasets or several years old; further, the quality of these datasets is generally poor, and they either do not create a test or benchmark dataset, or those datasets are randomly generated leading to leakage from the training data.
2.2 Minimizing data leakage
Most approaches to splitting a dataset into a training and a test set assume the data to be i.i.d. and perform a random split. A common exception is time-series data, where “future” data points are usually used for testing. However, data leakage kaufman2012leakage can still occur due to other factors. Kapoor et al. kapoor2022leakage provide a taxonomy of data leakage in machine learning, including cases where training and test data are not independent. A typical example is when multiple data points are collected from the same entity, such as a patient oner2020training , where random splitting can place data from the same entity in both sets, causing leakage. When entities are easily identifiable, data can be split to avoid overlap. In phishing detection, one potential source of leakage is sites built using the same kit. However, since such information is usually not available, we adopt a simple approach: we directly compare test and training points for similarity, dropping test points that are too similar to any training point. To make this search scalable to a large dataset, we use locality sensitive hashing andoni2008near .
2.3 Phishing detection approaches
One useful way to categorize phishing detection systems is by their deployment mode: (1) real-time, interactive detection – that blocks users from accessing harmful sites; and (2) offline detection – for generating blacklists and whitelists. The former imposes strict latency constraints, making computationally intensive methods or those requiring features that are slow to retrieve unsuitable for use. Our dataset is primarily designed to support the real-time detection scenario.
@@ 2.3 Phishing detection approaches
Phishing detection approaches have evolved over time tang2021survey ; zieni2023phishing ; li2024state , with early ones based on blacklists and whitelists of URLs. Although fast and effective for known sites, this approach suffers from complexity of maintaining these dynamic lists and its inability to detect zero-day phishing attacks. Subsequent methods used rules and heuristics based on features, which were later leveraged for machine learning models. The features can be broadly categorized into three types: (1) URL-based, (2) HTML-based, and (3) host-based.
@@
URL-based features refer to lexical features derived from the entire URL or its components ma2009beyond ; ma2011learning . Since URLs contain strong discriminative signals and are relatively short, a large number of datasets and methods focus solely on URLs-based features for detection el2020depth . HTML-based features are extracted from the HTML of a page, including its content, structure (DOM tree), tags, embedded links, and its visual rendering mohammad2015phishing . Host-based features relate to where a website is hosted and include its IP address, domain name properties, geographic information, and similar attributesma2009beyond . Due to their retrieval latency, these features are less suitable for real-time detection scenarios, which typically require response times under a few hundred milliseconds. Since our dataset is focused more towards real-time detection, it excludes host-based features.
@@
A wide range of machine learning approaches have been applied to phishing detection, spanning from hand-crafted features to learning representations of raw inputs. While many methods aim to directly classify phishing pages whittaker2010large , others have focused on identifying the impersonated brand first, usually using webpage snapshots and a reference dataset of known brands, followed by checking URL consistency afroz2011phishzoo ; abdelnabi2020visualphishnet ; lin2021phishpedia ; li2024phishintel . Most recently, LLMs are being used for detecting phishing pages li2024knowphish ; koide2024chatphishdetector ; cao2025phishagent ; zhang2025benchmarking . LLMs now have large enough context sizes that an entire HTML page can fit. Although they are computationally expensive, large datasets or model training is not required.
3 PhreshPhish Dataset
@@
We collected our dataset over eight months (July 2024 to March 2025), encompassing a broad range of phishing and benign URLs. Phishing URLs were primarily sourced from PhishTank phishtank , the AntiPhishing Working Group (APWG) eCrime eXchange apwgecrimex , and NetCraft netcraft . Benign URLs were drawn from anonymized browsing telemetry from over six million global Webroot users and Google search results for heavily targeted brands. These user-sourced URLs provide a more realistic and representative sample of benign pages that users are likely to encounter on the web as opposed to those from other datasets such as Common Crawl commoncrawl . The dataset was cleaned and curated to ensure that it is of high quality and free from noise and other issues that are common in other datasets.
3.1 Scraping challenges
Despite substantial research in web crawling, gathering data for applications in security remains difficult, particularly for phishing detection, due to its evasive and transient nature. Below, we outline key challenges and our mitigation strategies.
@@ 3.1 Scraping challenges
Table 1: Common failure modes encountered when scraping phishing webpages.
@@
Dynamic pages
Modern websites frequently rely on JavaScript frameworks like React react and Angular angular to load content dynamically. This poses a challenge for traditional crawlers, which only fetch static HTML and cannot execute client-side scripts. Consequently, dynamic content is often missed, limiting the usefulness of datasets such as Common Crawl and datasets derived from it raffel2020exploring ; penedo2024fineweb ; datacommons , which rely on traditional scraping tools like Apache Nutch nutch .
Ephemerality
Phishing pages are often short-lived and may be taken down before they can be scraped. This is particularly true for pages that are hosted on compromised benign domains, which are often taken down by the domain owner or hosting provider. This can lead to situations where the scraper is unable to access the original phishing page and instead is presented with a benign page or a takedown notice. This is a common failure mode that is frequently encountered when scraping phishing pages in particular.
Cloaking
Cloaking 9519414 is a technique used by attackers to present different content conditioned on specific criteria, such as a client’s IP address or user agent. This can lead to situations where the scraper is presented with a benign page instead of the phishing page, resulting in mislabeled data. This is a particularly pernicious problem that can lead to model poisoning tian2022comprehensive .
Staleness
Phishing techniques and tactics constantly evolve, with new pages and campaigns emerging regularly. In addition, underlying web technologies change over time, with new standards like HTML5 and frameworks like GraphQL graphql being adopted. Datasets must be regularly updated to reflect the current state of the web and of modern phishing techniques. Many existing phishing datasets are outdated and no longer relevant, as they were collected years ago and do not reflect the current state of the web. To combat the issue of staleness, we plan to maintain PhreshPhish with regular updates to ensure its relevance and utility for research purposes.
3.2 Our scraping process
To address the challenges outlined above, we developed a scraping pipeline that is designed to maximize the probability of capturing high-quality phishing data. Our scraping pipeline leverages Selenium selenium to render pages in full browser instances, capturing dynamic content missed by traditional tools. Pages are scraped using a distributed cluster running Windows and Chrome, allowing us to address the challenges of dynamic pages and cloaking. Nodes in the cluster are configured to use a variety of user agents and IP addresses to improve the chances of capturing the phishing page. To mitigate the ephemerality challenge, URLs are scraped within minutes of being reported to our feeds. Finally, we continuously monitor the performance of our scraping system and make adjustments as needed to ensure that it remains effective.
3.3 Cleaning process
Although our scraping process is designed to maximize the probability of capturing phishing data, bad data corresponding to one of the failure modes described in Table 1 is still common. After the raw data is collected, it undergoes a cleaning process to further improve data quality. This cleaning process is a combination of automated and manual steps where a number of heuristics are used to remove low-quality data.
Because URLs are often reported to multiple feeds, we normalize the URL to a canonical form and deduplicate the data at the URL level. The canonicalization process includes removing trailing slashes and converting the URL to lowercase 222 Although URLs can contain uppercase letters, we convert it to lowercase for deduplication purposes only since it is often the case that the same URL will be reported multiple times, occasionally with different casing. After deduplication, the original case-sensitive URL is retained. . Next, the HTML is parsed and the title of the page is extracted. Pages with titles containing certain keywords such as "forbidden" or "not found" are removed as they are likely to be scrape failures.
Pages are then grouped together and a representative page, known as a prototype, is selected from each group for manual inspection. A human annotator inspects the prototypical page and decides whether to keep or remove the page from the dataset. The decision is then applied to all nearest neighbors of the prototype with respect to cosine similarity of the TF-IDF features of the page. This process continues until a specified budget (i.e. the maximum number of prototypes to inspect) is exhausted. To maximize the efficiency of the cleaning process, we use two distinct grouping schemes. The first grouping is done by performing locality-sensitive hashing (LSH) via random projection. The second grouping is done with respect to the page title. The groupings are then ordered from largest to smallest with the largest groups being inspected first. Although the cleaning stage requires manual human inspection and annotation, it is optimized to an extent due to the fact that the human annotator only needs to inspect a representative page from each group and any decision made about the prototypical page is applied to all nearest neighbors in bulk. This significantly reduces the amount of manual work required to clean the dataset. Furthermore, a budget is imposed on the number of prototypes to be inspected to ensure that the cleaning process is efficient and does not take an excessive amount of time. Pages likely to contain personally identifiable information (PII) are also removed from the dataset. Further details including pseudocode for cleaning and our PII removal process are provided in Appendix C.
3.4 Final dataset
The final dataset is comprised of approximately 253k benign and 119k phishing pages, totaling 372k data points in all. Further details and exploratory analysis on the final dataset are provided in Appendix C.
@@ 3.4 Final dataset
Figure 2: A one-month snapshot top phishing targets and their domains. The targeted brands follow a power law-like distribution, with a few brands being targeted much more frequently than others. The most common domains in phishing URLs are often benign domains that allow users to upload and host content such as GitHub and Blogspot.
4 PhreshPhish Test and Benchmark Datasets

### v2

@@ 1. Introduction
Phishing attacks are a pervasive and evolving threat that uses social engineering techniques to deceive millions of Internet users annually. According to the latest Phishing Activity Trends Report (apwg-report-2025q1) released by the Anti-Phishing Working Group (APWG) (apwg), almost a million phishing attacks were observed just in the first quarter of 2025. These attacks are a central component of modern cyber kill chains (cyberkillchain), often involving campaigns that lure users to spoofed websites designed to trick them into revealing sensitive information, such as login credentials and financial data. The data thus obtained often serves as an initial entry point to fraud, identity theft and large-scale data breaches, resulting in substantial financial and reputational damage. In 2024 alone, phishing contributed significantly to the estimated \$16.6 billion in cyberfraud losses reported to the FBI (fbi-ic3report-2024).
@@
Phishing remains a popular attack vector due to its effectiveness, low cost, and ease of deployment. A major contributing factor is the widespread availability of phishing kits (cova2008there; bijmans2021catching). These kits contain complete phishing websites in a ready-to-deploy package, and can often be downloaded for free or even purchased on the dark web markets (dark-web-2018), making such attacks accessible even to unskilled adversaries.
@@
Despite considerable research efforts, detecting phishing sites remains a challenging task (das2019sok). Various approaches have been proposed to detect phishing, including use of blacklists, whitelists, heuristic-based techniques and machine learning methods. However, its adversarial nature, with attackers constantly adapting to avoid detection, and lack of good quality data undermines these defense approaches. Phishing websites are typically short-lived, highly dynamic, and often employ techniques such as obfuscation, cloaking (9519414), and fast-flux hosting (nagunwa2022machine).
Deployed phishing detection systems can be integrated into various applications, with browser plugins serving as a convenient option for verifying websites, and blocking phishing sites before users access them. To be effective in real-world settings, such systems must satisfy two key desiderata — (1) Low number of false postives: legitimate websites should only be rarely, if ever, blocked, implying a high precision, P(W|A)P(W|A) [^1] WW indicates a true phishing website; $A$ is a phishing website identified by a model., where 1P(W|A)1-P(W|A) represents the proportion of incorrectly blocked sites; and (2) Low latency: detection must operate in near real time, ideally within a few hundred milliseconds, to minimize user disruption.
@@
Figure 1. Our end-to-end pipeline consists of four distinct stages and we show the benign-to-phishing count at each step: (1) benign and phishing HTML is collected from the web using a real browser to ensure high fidelity; (2) the retrieved HTML is cleaned and assessed for quality using a combination of automated heuristics and human annotation; (3) train and test splits are created temporally and pruned to minimize leakage; (4) benchmark datasets are created by applying a set of diversity, difficulty and base rate filters.
@@
State-of-the-art machine learning-based phishing detection systems can be categorized into two broad approaches: (1) Direct phishing detection, where models are trained to classify webpages directly as phishing or benign (whittaker2010large; mohammad2015phishing; shahrivari2020phishing; aljofey2022effective). These models rely on feature extraction or representation learning from a (partially) labeled dataset of webpages. (2) Target brand-based phishing detection, where models are trained to identify the brand a webpage attempts to mimic, and phishing is inferred from inconsistencies between the webpage’s URL and the legitimate URLs associated with the identified brand (afroz2011phishzoo; lin2021phishpedia; li2024phishintel). While the latter approach does not require a dataset of phishing webpages, it relies on a curated set of target brands, and typically incurs much higher detection latency.
@@
While the collection of large-scale web datasets is itself nontrivial, successfully scraping phishing pages, which behave adversarially to avoid detection, presents a unique set of challenges beyond those typically encountered in general web scraping. In addition to the fact that the modern web is increasingly dynamic, issues such as ephemerality, cloaking, and staleness significantly complicate the collection of high-quality datasets which may be used to train and evaluate phishing detection models. These challenges necessitate the development of robust methodologies for collecting, cleaning, and validating phishing datasets to ensure their quality and utility for research purposes. We believe that these obstacles, discussed in greater detail in Section 3.1, contribute to the overall lack of high-quality datasets in this area. Although a search for publicly available phishing datasets would appear to yield promising results, we observed substantial quality issues with many datasets (see Section 3.2). We also could not find any suitable benchmark datasets that facilitated a robust comparative evaluation of detection techniques despite the significant role that phishing plays in cybercrime.
@@
In this work, we present PhreshPhish, a large-scale, high-quality phishing website dataset collected over more than one year. To the best of our knowledge, this is the largest publicly available phishing website dataset and offers significantly higher quality compared to other such public datasets. Further, we create realistic benchmark datasets for comparison between detection methods. Figure 1 provides a schematic of the entire data pipeline from obtaining the phishing and benign URLs to finally producing the training, test and benchmark datasets. To summarize, we make the following key contributions:
@@
  • We also release a suite of benchmark datasets to standardize the evaluation of detection methods. These benchmarks span five different base rates, ranging from 0.05% to 5%, making them more representative of real-world settings; further, we employ filters to increase dataset difficulty and diversity, and minimize data leakage.
@@
  • We evaluate and compare a number of baseline methods on our benchmark dataset to demonstrate the utility of our dataset and benchmarks. We also provide a detailed analysis of the performance of these methods.
2. Related Work
2.1. Phishing Detection
Phishing attacks span multiple channels, including email, SMS text messages (smishing), and social media (aleroud2017phishing). Detecting phishing at its point of origin (e.g., in email or SMS) can reduce exposure early in the attack lifecycle, but it typically requires channel-specific solutions and new channels such as voice (vishing) continue to emerge (enisa2025threatLandscape). A common downstream goal across channels is the harvesting of user credentials, which is often accomplished via a webpage. Accordingly, we focus our discussion on webpage-based phishing detection systems, which are commonly deployed either (1) in real-time to prevent users from accessing harmful sites, or (2) offline to support blacklist/whitelist generation. The former imposes strict latency constraints, making computationally intensive methods or those requiring features that are slow to retrieve unsuitable for use.
@@ 2.1. Phishing Detection
Webpage phishing detection approaches have evolved over time (tang2021survey; zieni2023phishing; li2024state), with early ones based on blacklists and whitelists of URLs. Although fast and effective for known sites, this approach suffers from complexity of maintaining these dynamic lists and its inability to detect zero-day phishing attacks. Subsequent methods used rules and heuristics based on features, which were later leveraged for machine learning models. The features can be broadly categorized into three types: (1) URL-based, (2) HTML-based, and (3) host-based.
@@
URL-based features refer to lexical features derived from the entire URL or its components (ma2009beyond; ma2011learning). Since URLs contain strong discriminative signals and are relatively short, a large number of datasets and methods focus solely on URLs-based features for detection (el2020depth). HTML-based features are extracted from the HTML of a page, including its content, structure (DOM tree), tags, embedded links, and its visual rendering (mohammad2015phishing). Host-based features relate to where a website is hosted and include its IP address, domain name properties, geographic information, and similar attributes(ma2009beyond). Due to their retrieval latency, these features are less suitable for real-time detection scenarios, which typically require response times under a few hundred milliseconds. Since our dataset is focused more towards real-time detection, it excludes host-based features.
@@
A wide range of machine learning approaches have been applied to phishing detection, spanning from hand-crafted features to learning representations of raw inputs. While many methods aim to directly classify phishing pages (whittaker2010large), others have focused on identifying the impersonated brand first, usually using webpage snapshots and a reference dataset of known brands, followed by checking URL consistency (afroz2011phishzoo; abdelnabi2020visualphishnet; lin2021phishpedia; li2024phishintel). Most recently, LLMs are being used for detecting phishing pages (li2024knowphish; koide2024chatphishdetector; cao2025phishagent; zhang2025benchmarking). LLMs now have large enough context sizes that an entire HTML page can fit. Although they are computationally expensive, large datasets or model training is not required.
2.2. Existing Datasets
Because many phishing detection systems are often focused on delivery mechanism (e.g., email, SMS, etc.), an internet search for phishing datasets appears to yield many results. However, when filtering for datasets suitable for training and evaluating webpage-based phishing detection systems, the results are much more limited. We searched for the terms "phishing" and "phish" on several common dataset hosting sites, including Hugging Face and Kaggle. We excluded smaller datasets – e.g., less than 10K rows on HuggingFace. Almost all datasets we found were focused on email-oriented phishing detection, with only a few datasets containing URL and/or HTML data.
Table 1. Phishing datasets on common hosting sites.
Repository Total Feature type
URL only Email/SMS URL & HTML
Hugging Face 35 15 14 2
Kaggle 19 6 4 4
Harvard Dataverse 2 0 2 0
OpenML 4 2 1 0
UCI ML Repo 3 3 0 0
Four datasets on Kaggle (kaggle-huntingdata11; kaggle-haozhang1579; kaggle-jackcavar; kaggle-aljofey) and two on Hugging Face (huggingface-ealvaradob; huggingface-huynq3Cyradar) include raw HTML, however, these datasets are either small, outdated, or exhibit substantial quality issues (see Section 3.2 for quality analysis). Additionally, we found most of the other datasets lacked standardized test or benchmark splits, and where splits were provided they were split randomly, which can lead to data leakage and over-optimistic performance estimates.
2.3. Minimizing Data Leakage
Most approaches to splitting a dataset into a training and a test set assume the data to be i.i.d. and perform a random split. A common exception is time-series data, where “future” data points are usually used for testing. However, data leakage (kaufman2012leakage) can still occur due to other factors. Kapoor et al. (kapoor2022leakage) provide a taxonomy of data leakage in machine learning, including cases where training and test data are not independent. A typical example is when multiple data points are collected from the same entity, such as a patient (oner2020training), where random splitting can place data from the same entity in both sets, causing leakage. When entities are easily identifiable, data can be split to avoid overlap. In phishing detection, one potential source of leakage is sites built using the same kit. However, since such information is usually not available, we adopt a simple approach: we directly compare test and training points for similarity, dropping test points that are too similar to any training point. To make this search scalable to a large dataset, we use locality sensitive hashing (andoni2008near).
3. PhreshPhish Dataset
@@
We collected our dataset over 17 months (July 2024–December 2025), encompassing a broad range of phishing and benign URLs. Phishing URLs were primarily sourced from PhishTank (phishtank), the AntiPhishing Working Group (APWG) eCrime eXchange (apwgecrimex), and NetCraft (netcraft). Benign URLs were drawn from anonymized browsing telemetry from over six million global Webroot users and Google search results for heavily targeted brands. These user-sourced URLs provide a more realistic and representative sample of benign pages that users are likely to encounter in practice, relative to benign URLs drawn from web-crawl datasets such as Common Crawl (commoncrawl). The dataset was cleaned and curated to reduce common sources of noise and labeling artifacts observed in prior datasets.
3.1. Scraping Challenges
Web scraping is a standard component of many data collection pipelines designed to collect large training corpora for domains ranging from natural language processing to computer vision. However, data collection for applications in computer security remains difficult, particularly in the phishing domain, due to the adversarial nature of the setting. In this section we identify common challenges encountered when collecting phishing datasets and describe the mitigation strategies incorporated into our pipeline. We posit that these factors contribute to the scarcity of large, high-quality datasets in the phishing literature.
@@ 3.1. Scraping Challenges
Table 2. Common failure modes encountered when scraping phishing webpages.
@@
Figure 2. A one-month snapshot (August 2025) of collected phishing pages. (2(a)) Most frequently targeted brands over time. Targeted brands follow a power law-like distribution, with a few brands being targeted much more frequently than others. (2(b)) The most common domains in phishing URLs are often legitimate domains that allow users to upload and host content such as Vercel and Blogspot.
Dynamic pages Modern websites rely on JavaScript and frameworks like React (react) and Angular (angular) to load content dynamically. This poses a challenge for traditional crawlers, which only fetch static HTML and cannot execute client-side scripts. Consequently, dynamic content is often missed, limiting the usefulness of datasets such as Common Crawl and datasets derived from it (raffel2020exploring; penedo2024fineweb; datacommons), which rely on traditional scraping tools like Apache Nutch (nutch).
Ephemerality Phishing pages are often short-lived and may be taken down before they can be scraped. This is particularly true for pages that are hosted on compromised benign domains, which are often taken down by the domain owner or hosting provider. This can lead to situations where the scraper is unable to access the original phishing page and instead is presented with a benign page or a takedown notice. Datasets that do not address this challenge risk containing significant amounts of mislabeled data.
Cloaking Cloaking (9519414) is a technique used by attackers to present different content conditioned on specific criteria, such as a client’s IP address or user agent. This can lead to situations where the scraper is presented with a benign page instead of the phishing page, resulting in mislabeled data. This adversarial technique can lead to model poisoning (tian2022comprehensive) if the mislabeled data is used for training, and can lead to over-optimistic performance estimates if the mislabeled data is used for evaluation.
Staleness Phishing techniques and tactics constantly evolve, with new pages and campaigns emerging regularly. In addition, underlying web technologies change over time, with new standards like HTML5 and frameworks like GraphQL (graphql) being adopted. Although concept drift is a challenge for all machine learning applications, it is particularly acute in phishing detection where evasion tactics evolve rapidly. Many existing phishing datasets are outdated and no longer relevant, as they were collected years ago and do not reflect the current state of the web. To combat the issue of staleness, we plan to maintain PhreshPhish with regular updates to ensure its relevance and utility.
To address the challenges outlined above, we developed a scraping pipeline that is designed to maximize the probability of capturing high-quality phishing data. Our scraping pipeline leverages Selenium (selenium) to render pages in full browser instances, capturing dynamic content missed by traditional tools. Pages are scraped using a distributed cluster running Windows and Chrome, allowing us to address the challenges of dynamic pages and cloaking. To further defend against cloaking, nodes in the cluster are configured to use a variety of user agents and IP addresses. To mitigate the ephemerality challenge, URLs are scraped within minutes of being reported to our feeds, maximizing the likelihood of capturing the original phishing page before it is taken down. Finally, we continuously monitor the performance of our scraping system and make adjustments as needed to ensure that it remains effective.
3.2. Cleaning Process
Although our scraping system is designed to maximize the probability of capturing live phishing pages, scrape failures corresponding to the failure modes outlined in Section 3.1 can still occur. Therefore, we developed a cleaning process designed to remove such failures from the dataset. This cleaning process consists of two stages: a heuristics-based automated filtering stage and a human-in-the-loop manual inspection stage.
Automated filtering Because URLs are often reported to multiple feeds, we normalize the URL to a canonical form and deduplicate the data at the URL level. The canonicalization process includes removing trailing slashes and converting the URL to lowercase  [^2] Although a URL can contain uppercase letters, we convert it to lowercase only for deduplication purposes since it is often the case that the same URL will be reported multiple times, occasionally with different casing. After deduplication, the original case-sensitive URL is retained. . Next, the HTML is parsed and the title of the page is extracted. Pages with titles containing certain keywords such as "forbidden" or "not found" are removed as they are likely to be scrape failures.
Manual inspection After automated filtering, pages are then grouped together and a representative page, known as a prototype, is selected from each group for manual inspection. A human annotator inspects the prototypical page and decides whether to keep or remove the page from the dataset. The decision is then applied to all nearest neighbors of the prototype with respect to cosine similarity of the TF-IDF features of the page HTML. This process continues until a specified budget (i.e. the maximum number of prototypes to inspect) is exhausted. To maximize the efficiency of the cleaning process, we use two distinct grouping schemes. The first grouping is done by extracting TF-IDF features from the page HTML and performing locality-sensitive hashing (LSH) via random projection to group similar pages together. The second grouping is done with respect to page title. For each scheme, the groupings are then ordered from largest to smallest with larger groups being inspected first. Although the cleaning stage requires manual human inspection and annotation, it is optimized to an extent due to the fact that the human annotator only needs to inspect a representative page from each group and any decision made about the prototypical page is applied to all nearest neighbors in bulk. This significantly reduces the amount of manual work required to clean the dataset. Furthermore, a budget is imposed on the number of prototypes to be inspected to ensure that the cleaning process is efficient and does not take an excessive amount of time. In practice, this manual stage is lightweight; a typical annotation budget is 1–3 hours.
Finally, pages likely to contain personally identifiable information (PII) are also removed from the dataset. Full details including pseudocode for cleaning and our PII removal process are provided in Appendix A.
3.3. Quality Assessment
We next assess dataset quality by comparing PhreshPhish to two prominent datasets in the literature: Aljofey, et al. (aljofey2022effective) and Crawling2024 (xie2025scalable). We apply the same cleaning methodology used to construct our dataset (Section 3.2) and use the observed rejection rates as a proxy for dataset quality. Because complete manual inspection is not feasible, we allocate a manual cleaning budget of approximately one hour of annotator time per dataset (60 LSH bins and 50 title groups). Then we estimate a lower and upper bound for the amount of data remaining if cleaning were to continue to completion. The results of our quality assessment are summarized in Table 3.
Table 3. Our two-stage cleaning pipeline applied to three phishing datasets.
Aljofey, et al. Crawling2024 PhreshPhish (ours)
Phish Benign Total Phish Benign Total Phish Benign Total
Stage 0: Initial Size
Initial Size ($N$) 27,280 32,972 60,252 26,687 29,428 56,115 250,002 366,991 616,993
Stage 1: Automated Filtering
URL Duplicates 2,936 129 3,065 1,069 166 1,235 0 0 0
HTML Missing 1,843 1,103 2,946 1,427 103 1,530 0 0 0
Bad Title 12,715 5,018 17,733 6,322 665 6,987 0 0 0
Remaining after Stage 1 12,722 26,851 39,573 18,938 28,660 47,598 250,002 366,991 616,993
Stage 2: Manual Inspection
LSH Rejections 2,530 5,422 7,952 3,390 3,470 6,860 1,927 790 2,717
Title Rejections 794 183 977 555 47 602 491 0 491
Remaining after Stage 2 9,398 21,246 30,644 14,993 25,143 40,136 247,584 366,201 613,785
Total Removed 17,882 11,726 29,608 11,694 4,285 15,979 2,418 790 3,208
Rejection Rate ($r$) 65.5% 35.6% 49.1% 43.8% 14.6% 28.5% 0.97% 0.02% 0.52%
Aljofey, et al. (aljofey2022effective) This dataset contains 60k samples, with 27k labeled as phishing and 33k labeled benign. We found this dataset contained a large number of data points that could immediately be removed under our automated filtering process (the majority of which were due to having bad titles such as "404 Not Found"). After running our cleaning process, over 65% of the phishing samples in this dataset were removed, leaving only 9,398 phish samples remaining in the dataset. The benign samples were also heavily affected, with nearly 36% of the benign samples being removed. Assuming the remaining un-inspected samples are perfectly clean, the dataset would contain 9,398 phishing samples and 21,246 benign samples. Under more realistic assumptions and based on rejection rates observed, however, we estimate that the dataset would contain only around 3,242 phishing samples and 13,682 benign samples if cleaning were to continue to completion.
Crawling2024 (xie2025scalable) This dataset contains 56k samples, with 27k labeled as phishing and 29k labeled as benign. Although this dataset performed better than Aljofey, et al., there were still significant quality issues with over 43% of phishing samples being marked for removal during cleaning. The dataset would contain 14,993 phishing samples and 25,143 benign samples for a total of 40,136 samples as a upper bound. Under more realistic assumptions and based on rejection rates observed, however, we estimate that the dataset would contain only around 8,400 phishing samples and 20,000 benign samples if cleaning were to continue to completion.
PhreshPhish (ours) Combining the train and full benchmark set, we have a total of 617k samples, with 250k labeled as phishing and 367k labeled as benign. Because we already previously applied our cleaning process to the dataset, the automated filtering stage didn’t result in any new samples being removed. Upon manual inspection, fewer than 1% of the phishing samples and a negligible number of benign samples were removed. This is expected, and we do not claim that our dataset is free of all errors, as an unlimited cleaning budget is not feasible. Nevertheless, the low rejection rate provides evidence of higher dataset quality relative to prior collections under the same auditing protocol. Moreover, PhreshPhish is substantially larger than existing datasets, with nearly 248k phishing samples and 366k benign samples remaining after cleaning (approximately an order of magnitude larger than the other datasets).
4. Test and Benchmark Datasets